OpenStudy (anonymous):

you draw a card from a standard deck of 52 cards. Find the probability of drawing a face card

OpenStudy (anonymous):

how many face cards there are in the deck divided by the total number of cards in the deck

OpenStudy (anonymous):

ok so it would be... 12 divide by 52?

OpenStudy (anonymous):

yes, after you reduce

OpenStudy (anonymous):

\[\frac{12}{52}=\frac{3}{13}\] so you can think if it as the number of face cards in each suit

OpenStudy (anonymous):

so it would be 0.230769231?

OpenStudy (anonymous):

i would leave it as \(\frac{3}{13}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw btw you decimal is right, so if you need the decimal form you are good with that

OpenStudy (anonymous):

kk